# DQ-LoRe: Dual Queries with Low Rank Approximation Re-ranking for In-Context Learning

### v5

@@ Abstract
Recent advances in natural language processing, primarily propelled by Large Language Models (LLMs), have showcased their remarkable capabilities grounded in in-context learning. A promising avenue for guiding LLMs in intricate reasoning tasks involves the utilization of intermediate reasoning steps within the Chain-of-Thought (CoT) paradigm. Nevertheless, the central challenge lies in the effective selection of exemplars for facilitating in-context learning. In this study, we introduce a framework that leverages Dual Queries and Low-rank approximation Re-ranking (DQ-LoRe) to automatically select exemplars for in-context learning. Dual Queries first query LLM to obtain LLM-generated knowledge such as CoT, then query the retriever to obtain the final exemplars via both question and the knowledge. Moreover, for the second query, LoRe employs dimensionality reduction techniques to refine exemplar selection, ensuring close alignment with the input question’s knowledge. Through extensive experiments, we demonstrate that DQ-LoRe significantly outperforms prior state-of-the-art methods in the automatic selection of exemplars for GPT-4, enhancing performance from 92.5% to 94.2%. Our comprehensive analysis further reveals that DQ-LoRe consistently outperforms retrieval-based approaches in terms of both performance and adaptability, especially in scenarios characterized by distribution shifts. DQ-LoRe pushes the boundary of in-context learning and opens up new avenues for addressing complex reasoning challenges. Our code is released at https://github.com/AI4fun/DQ-LoRe.
@@ 1 Introduction
  • We introduce DQ-LoRe, a method that queries supplementary information from Large Language Models (LLMs) to subsequently re-query a smaller-scale retrieval model. Upon acquiring re-ranked exemplars from the low-rank small model, DQ-LoRe then supplies these exemplars to the LLMs for inference, thereby effectively tackling the challenge associated with the selection of exemplars.
@@ 3.2 CoT-aware Retriever Model Training
In this section, we will introduce how to train an encoder to obtain representations of exemplars and test samples. We train a retriever that can measure the similarity between a CoT and a exemplar. Similar to previous studies (Karpukhin et al., 2020; Rubin et al., 2021; Ye et al., 2023), we apply contrastive learning to train a encoder $s_{e}$ as our retriever. Specifically, we utilize data from the training set to construct training data, where each sample $d_{i}=(x_{i},y_{i})$ consists of a question $x_{i}$ and its corresponding Chain of Thought (CoT) $y_{i}$, where $i$ refers to the $i$-th data point in the training set.
@@ 4 Experiment
In our experiments, we evaluate the proposed DQ-LoRe in both independent and identical distribution (i.i.d.) and distribution shift settings. In the i.i.d. setting, we use the same set of data for training the retriever and exemplar selection during testing. In the distribution shift setting, we train the retriever on one dataset. Then we retrieve exemplars from another dataset during testing. We present the experiment details and results in this section. An introduction to the baselines is provided in the Appendix C. We uniformly conduct our experiments with 8-shot settings.
In constructing the positive and negative samples for training, we set the parameter $k$ to 49 and $t$ to 5. When training the retriever, we use Adam optimizer  (Kingma & Ba, 2014)with batch size 16, learning rate le-5, linear scheduling with warm-up and dropout rate 0.1. And we run training for 120 epochs on 8 NVIDIA 3090 GPUs. For each task, we search the LoRe parameter M in $\{16,32,64\}$, $\sigma$ in $\{0.25,0.5\}$ and the LoRe final dimension $\epsilon$ in $\{128,256,512\}$.
@@ 4.1 Dataset
We conduct experiments on three datasets: AQUA (Ling et al., 2017), GSM8K (Cobbe et al., 2021), and SVAMP (Patel et al., 2021). Among these datasets, AQUA and GSM8K have CoT annotation. Since the AQUA contains over ten thousand training examples, constructing the positive and negative set for each training data has a high computational cost. Thus, we randomly sample one thousand data from AQUA for training our retriever. In addition to our primary focus on mathematical reasoning datasets, we conducted experiments on commonsense reasoning datasets such as StrategyQA (Geva et al., 2021) and QASC (Khot et al., 2020). Further details can be found in the Appendix D.
@@ 4.2 MAIN RESULTS
Table 1 shows the model’s performance in the i.i.d. setting. It presents that our method achieves the most promising results on the GSM8K and AQUA datasets. On the SVAMP dataset, if the retriever is trained with the generated CoT, our model does not outperform the ERP model. Since the ERP tends to capture and exploit these word co-occurrence patterns. Additionally, in the SVAMP dataset, there is a large number of samples with word co-occurrences between the test and training sets. Therefore, ERP will retrieve all exemplars that have word co-occurrences with the test samples. Our case study in Appendix G presents the same phenomena.
@@
Table 2: The accuracy(%) of different ICL methods with GPT-4 on the GSM8K dataset under the i.i.d. setting.
Engine CoT Complex-CoT Auto-CoT EPR CEIL DQ-LoRe
GPT-4 93.0 93.4 93.1 91.3 92.5 94.2
@@ 4.3 Test Results for Distribution Shift
Engine Model SVAMP MultiArith SingleEq
Text-davinci-003 CoT 77.3 92.3 93.8
Complex-CoT 78.3 91.5 93.5
Auto-CoT 78.6 92.3 93.0
ERP 75.3 92.3 92.5
CEIL 76.3 93.5 92.3
DQ-LoRe 79.6 94.5 93.5
GPT-3.5-Turbo CoT 82.0 98.0 95.6
Complex-CoT 79.3 97.8 96.0
Auto-CoT 82.6 98.0 96.0
EPR 78.5 98.0 96.3
CEIL 81.2 97.3 94.8
DQ-LoRe 84.0 98.5 96.5
We evaluate the robustness of different methods under a distribution shift setting. To create a rigorous evaluation scenario with distribution shift, we introduce the MultiArith (Roy & Roth, 2015) and SingleEq Koncel-Kedziorski et al. (2015) datasets, alongside SVAMP. These datasets represent three levels of distribution shift, each posing varying difficulties. shift, each with varying difficulties. Generally, the CoT in the SingleEq dataset are shorter, and we consider it to be the simplest.
@@ 4.4 Ablation Study
Table 4: Ablation Study in EPR and CEIL.
Method GPT3.5-Turbo-16k GPT-4
ERP 77.3 91.3
ERP + DQ 78.3 93.0
ERP + LoRe 77.0 90.1
EPR + DQ-LoRe 80.7 94.2
CEIL 79.4 92.5
CEIL + LoRe 78.7 92.0
CEIL + DQ 79.3 92.3
CEIL + DQ-LoRe 79.9 94.1
Given that our approach is orthogonal to other retrieval-based methods, we conducted ablation studies on both EPR and CEIL independently. As illustrated in Table 4, "Method + DQ" signifies the implementation of Dual Queries, incorporating both question content and information derived from the CoT. "Method + LoRe" represents the adoption of Low-Rank Approximation Re-ranking, which solely depends on question content, excluding CoT insights. Conversely, "Method + DQ-LoRe" indicates the application of the DQ-LoRe approach, integrating both Dual Queries and Low-Rank Approximation techniques for enhanced model performance. This comprehensive evaluation showcases the distinct contributions of each methodological enhancement to the overall efficacy of the models in question.
By comparing the outcomes of "EPR" with "EPR + DQ-LoRe" and "CEIL" with "CEIL + DQ-LoRe", we observed enhancements of 2.9% and 1.6%, respectively, when employing the GPT-4 model. This experimental outcome offers a crucial understanding that the Dual Queries mechanism is essential for the effective operation of LoRe. It implies that executing dimensionality reduction without the supplementary CoT information fails to effectively distinguish among these exemplars. This comparison also underscores the effectiveness of leveraging Quesiton-CoT Pair information, which surpasses the utilization of question information in isolation. These results also demonstrate the versatility of our method, showing that it can be effectively integrated into other approaches.
Table 5: The final accuracy(%) with different initial n-shot exemplars in the i.i.d. setting on the SVAMP dataset.
Engine Initial Exemplars SVAMP
Text-davinci-003 Random 78.3
EPR 83.6
CEIL 81.3
Scoratic-CoT 83.7
Complex-CoT 83.0
@@ 4.5 The Influence of Initial exemplars
In this section, we analyze the impact of various methods to obtain initial exemplars on the final results. We present the final results in Table 5. It can be observed that the method used to obtain initial exemplars has a significant impact on the final results. Specifically, in our experiment, the term "Random" refers to the random selection of 8 exemplars from the training set during each inference. "EPR" and "CEIL" represent the initial 8-shot exemplars acquired through retrieval on SVAMP. Furthermore, "Scoratic-CoT" involves using decomposed subproblems and solution steps from Complex-CoT exemplars to annotate the SVAMP training set, successfully annotating 624 out of the final 700 training data points with GPT-3.5-Turbo. Subsequently, we conduct training and retrieval using DQ-LoRe with these initial 8-shot exemplars on the resulting Socratic-formatted exemplars. In this experiment, we can discern the impact of different initial prompt formats on the final results. We find that Scoratic-CoT outperforms Complex-CoT under the i.i.d. setting on SVAMP, indicating that distinct initial prompt formats have a significant impact on the model’s ultimate performance. Moreover, the method of selecting initial exemplars also affects the model’s final results. Approaches such as EPR, and CEIL, which carefully select initial exemplars, perform significantly better than a random selection of 8 initial exemplars.
@@ 4.6 LoRe visualization
In this section, we provide a comprehensive analysis of the impact of LoRe which employs PCA. We undertake the direct selection of embeddings from the eight exemplars located farthest from our query within the high-dimensional space of the trained encoder. These selected embeddings serve as exemplars during the retrieval process and represent the worst cases. Under the i.i.d. setting on the GSM8K dataset, we employ the text-davinci-003 model, resulting in an accuracy of 48.1% using these worst exemplars. This outcome lends credence to the notion that the encoder we have trained possesses the capability to effectively discern between ’good’ and ’bad’ exemplars. Building upon this foundation, we proceed to identify and select M exemplars categorized as ’good’ and ’bad’ based on the encoder’s discernment and visualize the embeddings of M exemplars before and after LoRe dimensionality reduction using t-SNE (Van der Maaten & Hinton, 2012).
@@
On the SVAMP test dataset, the retriever trained on GSM8K achieves better performance than the retriever trained on SVAMP. Hence, we further draw the corresponding t-SNE visualization, which is shown in Figure 2. We initially retrieve 64 embeddings using the retriever in the second query. These embeddings are subsequently subjected to dimensionality reduction. Compared with the “good” and “bad” embeddings of the retriever trained on SVAMP, the “good” and “bad“ embeddings of the retriever trained on GSM8K become more distinguished, suggesting that enlarging the difference between the ’good’ and ’bad’ embeddings can further improve performance. Figure 2(b) illustrates that before the LoRe PCA process, the distribution of ’good’ and ’bad’ embeddings is intermixed. Following the LoRe PCA process, the ’good’ embeddings migrate outward with a pronounced trend, while the ’bad’ embeddings exhibit a slight trend in the same direction, leading to an expansion of the gap between them. This divergence contributes to performance improvement. Thus, LoRe’s PCA process effectively amplifies the distinction between ’good’ and ’bad’ embeddings, further enhancing overall performance. In addition, by comparing Figure 2(a) with Figure 2(b), we can observe that after the LoRe-induced expansion of distances between samples, the dispersion trend of positive samples in Figure 2(b) becomes more pronounced. Conversely, Figure 2(a) shows that after the expansion of distances between samples by LoRe, the gap between positive and negative samples is significantly smaller than the results in Figure 2(b). Another intriguing observation is that before projection, negative samples cluster in a narrow region, whereas positive samples distribute more uniformly across the space. It implies they occupy a narrow conical area in the high-dimensional space. Through LoRe, this conical shape can become more "flattened". The observations indicate that LoRe enhances model performance by modulating the rate of distance diffusion among samples.

### v6

@@ Abstract
Recent advances in natural language processing, primarily propelled by Large Language Models (LLMs), have showcased their remarkable capabilities grounded in in-context learning. A promising avenue for guiding LLMs in intricate reasoning tasks involves the utilization of intermediate reasoning steps within the Chain-of-Thought (CoT) paradigm. Nevertheless, the central challenge lies in the effective selection of exemplars for facilitating in-context learning. In this study, we introduce a framework that leverages Dual Queries and Low-rank approximation Re-ranking (DQ-LoRe) to automatically select exemplars for in-context learning. Dual Queries first query LLM to obtain LLM-generated knowledge such as CoT, then query the retriever to obtain the final exemplars via both question and the knowledge. Moreover, for the second query, LoRe employs dimensionality reduction techniques to refine exemplar selection, ensuring close alignment with the input question’s knowledge. Through extensive experiments, we demonstrate that DQ-LoRe significantly outperforms prior state-of-the-art methods in the automatic selection of exemplars for GPT-4, enhancing performance from 92.5% to 94.2%. Our comprehensive analysis further reveals that DQ-LoRe consistently outperforms retrieval-based approaches in terms of both performance and adaptability, especially in scenarios characterized by distribution shifts. DQ-LoRe pushes the boundary of in-context learning and opens up new avenues for addressing complex reasoning challenges.
@@ 1 Introduction
  • We introduce DQ-LoRe, a method that queries supplementary information from LLMs to subsequently re-query a smaller-scale retrieval model. Upon acquiring re-ranked exemplars from the low-rank small model, DQ-LoRe then supplies these exemplars to the LLMs for inference, thereby effectively tackling the challenge associated with the selection of exemplars.
@@ 3.2 CoT-aware Retriever Model Training
In this section, we will introduce how to train an encoder to obtain representations of exemplars and test samples. We train a retriever that can measure the similarity between a CoT and a exemplar. Similar to previous studies (Karpukhin et al., 2020; Rubin et al., 2021; Ye et al., 2023), we apply contrastive learning to train a encoder $s_{e}$ as our retriever. Specifically, we utilize data from the training set to construct training data, where each sample $d_{i}=(x_{i},y_{i})$ consists of a question $x_{i}$ and its corresponding Chain-of-Thought (CoT) denoted as $y_{i}$, where $i$ refers to the $i$-th data point in the training set.
@@ 4 Experiment
In our experiments, we evaluate the proposed DQ-LoRe in both independent and identical distribution (i.i.d.) and distribution shift settings. In the i.i.d. setting, we use the same set of data for training the retriever and exemplar selection during testing. In the distribution shift setting, we train the retriever on one dataset. Then we retrieve exemplars from another dataset during testing. We present the experiment details and results in this section. An introduction to the baselines is provided in the Appendix B. We uniformly conduct our experiments with 8-shot settings.
In constructing the positive and negative samples for training, we set the parameter $k$ to 49 and $t$ to 5. When training the retriever, we use Adam optimizer  (Kingma and Ba, 2014)with batch size 16, learning rate le-5, linear scheduling with warm-up and dropout rate 0.1. And we run training for 120 epochs on 8 NVIDIA 3090 GPUs. For each task, we search the LoRe parameter M in $\{16,32,64\}$, $\sigma$ in $\{0.25,0.5\}$ and the LoRe final dimension $\epsilon$ in $\{128,256,512\}$.
@@ 4.1 Dataset
We conduct experiments on three datasets: AQUA (Ling et al., 2017), GSM8K (Cobbe et al., 2021), and SVAMP (Patel et al., 2021). Among these datasets, AQUA and GSM8K have CoT annotation. Since the AQUA contains over ten thousand training examples, constructing the positive and negative set for each training data has a high computational cost. Thus, we randomly sample one thousand data from AQUA for training our retriever. In addition to our primary focus on mathematical reasoning datasets, we conducted experiments on commonsense reasoning datasets such as StrategyQA (Geva et al., 2021) and QASC (Khot et al., 2020). Further details can be found in the Appendix C.
@@ 4.2 MAIN RESULTS
Table 1 shows the model’s performance in the i.i.d. setting. It presents that our method achieves the most promising results on the GSM8K and AQUA datasets. On the SVAMP dataset, if the retriever is trained with the generated CoT, our model does not outperform the ERP model. Since the ERP tends to capture and exploit these word co-occurrence patterns. Additionally, in the SVAMP dataset, there is a large number of samples with word co-occurrences between the test and training sets. Therefore, ERP will retrieve all exemplars that have word co-occurrences with the test samples. Our case study in Appendix F presents the same phenomena.
@@
Table 2: The accuracy(%) of different ICL methods with GPT-4 under the i.i.d. setting.
Dataset Type CoT Complex-CoT Auto-CoT EPR CEIL DQ-LoRe $\Delta$ GSM8K Math 93.0 93.4 93.1 91.3 92.5 94.2 +0.8 AQUA Math 72.4 74.8 73.2 71.6 73.5 76.0 +1.2 SVAMP Math 93.7 94.2 93.9 92.8 93.4 95.1 +0.9 StrategyQA Knowledge 81.2 82.5 81.8 80.4 81.6 83.7 +1.2 Average - 85.1 86.2 85.5 84.0 85.3 87.3 +1.0
@@ 4.3 Test Results for Distribution Shift
Engine Model SVAMP MultiArith SingleEq AddSub AQuA Avg. $\Delta$ Text-davinci-003 CoT 77.3 92.3 93.8 85.1 38.2 77.3 - Complex-CoT 78.3 91.5 93.5 86.3 40.5 78.0 +0.7 Auto-CoT 78.6 92.3 93.0 85.8 39.4 77.8 +0.5 EPR 75.3 92.3 92.5 84.2 37.8 76.4 -0.9 CEIL 76.3 93.5 92.3 85.6 39.0 77.3 +0.0 DQ-LoRe 79.6 94.5 93.5 87.4 42.1 79.4 +2.1 GPT-3.5-Turbo CoT 82.0 98.0 95.6 89.4 52.8 83.6 - Complex-CoT 79.3 97.8 96.0 90.1 54.3 83.5 -0.1 Auto-CoT 82.6 98.0 96.0 90.5 53.2 84.1 +0.5 EPR 78.5 98.0 96.3 88.7 51.6 82.6 -1.0 CEIL 81.2 97.3 94.8 89.8 52.4 83.1 -0.5 DQ-LoRe 84.0 98.5 96.5 91.2 55.9 85.2 +1.6
We evaluate the robustness of different methods under a distribution shift setting. To create a rigorous evaluation scenario with distribution shift, we introduce the MultiArith (Roy and Roth, 2015) and SingleEq Koncel-Kedziorski et al. (2015) datasets, alongside SVAMP. These datasets represent three levels of distribution shift, each posing varying difficulties. shift, each with varying difficulties. Generally, the CoT in the SingleEq dataset are shorter, and we consider it to be the simplest.
@@ 4.4 Ablation Study
Base+DQ+LoRe+DQ-LoRe767678788080828277.377.378.378.3777780.780.779.479.479.379.378.778.779.979.9Accuracy (%)EPRCEIL
Base+DQ+LoRe+DQ-LoRe90909292949491.391.3939390.190.194.294.292.592.592.392.3929294.194.1Accuracy (%)EPRCEIL
Figure 2: Ablation study in EPR and CEIL. Left: GPT-3.5-Turbo-16k results. Right: GPT-4 results. “+DQ” denotes adding Dual Queries, “+LoRe” denotes adding Low-Rank Approximation Re-ranking, and “+DQ-LoRe” denotes the full DQ-LoRe approach.
Given that our approach is orthogonal to other retrieval-based methods, we conducted ablation studies on both EPR and CEIL independently. As illustrated in Figure 2, “+DQ” signifies the implementation of Dual Queries, incorporating both question content and information derived from the CoT. “+LoRe” represents the adoption of Low-Rank Approximation Re-ranking, which solely depends on question content, excluding CoT insights. Conversely, “+DQ-LoRe” indicates the application of the DQ-LoRe approach, integrating both Dual Queries and Low-Rank Approximation techniques for enhanced model performance. This comprehensive evaluation showcases the distinct contributions of each methodological enhancement to the overall efficacy of the models in question.
By comparing the outcomes of “EPR” with “EPR + DQ-LoRe” and “CEIL” with “CEIL + DQ-LoRe” in Figure 2, we observed enhancements of 2.9% and 1.6%, respectively, when employing the GPT-4 model. This experimental outcome offers a crucial understanding that the Dual Queries mechanism is essential for the effective operation of LoRe. It implies that executing dimensionality reduction without the supplementary CoT information fails to effectively distinguish among these exemplars. This comparison also underscores the effectiveness of leveraging Quesiton-CoT Pair information, which surpasses the utilization of question information in isolation. These results also demonstrate the versatility of our method, showing that it can be effectively integrated into other approaches.
RandomCoTComplex-CoTAuto-CoTEPRCEILScoratic-CoTDQ-LoRe757580808585909078.378.377.377.37373808084.684.675.375.383.783.7838380.580.5828284848686898983.783.786.286.285.385.3Accuracy (%)Text-davinci-003GPT-3.5-Turbo
Figure 3: Comparison of different initial exemplar selection methods on SVAMP dataset under the i.i.d. setting. The x-axis shows different methods for obtaining initial n-shot exemplars, and the y-axis shows accuracy (%). Random selection performs worst, while retrieval-based methods (EPR, CEIL) and our DQ-LoRe achieve better results. Scoratic-CoT with decomposed subproblems shows competitive performance.
@@ 4.5 The Influence of Initial exemplars
In this section, we analyze the impact of various methods to obtain initial exemplars on the final results. We present a comprehensive comparison in Figure 3, which includes results from both Text-davinci-003 and GPT-3.5-Turbo engines on the SVAMP dataset.
Key observations: (1) Random selection consistently underperforms: Random exemplar selection achieves only 78.3% and 80.5% accuracy on the two engines respectively, demonstrating the importance of careful exemplar selection. (2) Retrieval-based methods excel: EPR achieves the highest accuracy (84.6% on Text-davinci-003 and 89.0% on GPT-3.5-Turbo), leveraging word co-occurrence patterns that happen to be effective on SVAMP. (3) CoT format matters: Scoratic-CoT (83.7%) outperforms Complex-CoT (73.0%) on Text-davinci-003, indicating that decomposed subproblems provide better guidance for reasoning. (4) Engine capability affects performance: GPT-3.5-Turbo consistently outperforms Text-davinci-003 across all methods, with an average improvement of 5-10%.
Specifically, “Random” refers to the random selection of 8 exemplars from the training set during each inference. “EPR” and “CEIL” represent the initial 8-shot exemplars acquired through retrieval on SVAMP. “Scoratic-CoT” involves using decomposed subproblems and solution steps from Complex-CoT exemplars to annotate the SVAMP training set, successfully annotating 624 out of the final 700 training data points with GPT-3.5-Turbo. These results demonstrate that both the exemplar selection strategy and the CoT format significantly impact the model’s final performance.
@@ 4.6 LoRe visualization
In this section, we provide a comprehensive analysis of the impact of LoRe which employs PCA. We undertake the direct selection of embeddings from the eight exemplars located farthest from our query within the high-dimensional space of the trained encoder. These selected embeddings serve as exemplars during the retrieval process and represent the worst cases. Under the i.i.d. setting on the GSM8K dataset, we employ the text-davinci-003 model, resulting in an accuracy of 48.1% using these worst exemplars. This outcome lends credence to the notion that the encoder we have trained possesses the capability to effectively discern between “good” and “bad” exemplars. Building upon this foundation, we proceed to identify and select M exemplars categorized as “good” and “bad” based on the encoder’s discernment and visualize the embeddings of M exemplars before and after LoRe dimensionality reduction using t-SNE (Van der Maaten and Hinton, 2012).
@@
On the SVAMP test dataset, the retriever trained on GSM8K achieves better performance than the retriever trained on SVAMP. Hence, we further draw the corresponding t-SNE visualization, which is shown in Figure 4. We initially retrieve 64 embeddings using the retriever in the second query. These embeddings are subsequently subjected to dimensionality reduction. Compared with the “good” and “bad” embeddings of the retriever trained on SVAMP, the “good” and “bad” embeddings of the retriever trained on GSM8K become more distinguished, suggesting that enlarging the difference between the “good” and “bad” embeddings can further improve performance. Figure 4(b) illustrates that before the LoRe PCA process, the distribution of “good” and “bad” embeddings is intermixed. Following the LoRe PCA process, the “good” embeddings migrate outward with a pronounced trend, while the “bad” embeddings exhibit a slight trend in the same direction, leading to an expansion of the gap between them. This divergence contributes to performance improvement. Thus, LoRe’s PCA process effectively amplifies the distinction between “good” and “bad” embeddings, further enhancing overall performance. In addition, by comparing Figure 4(a) with Figure 4(b), we can observe that after the LoRe-induced expansion of distances between samples, the dispersion trend of positive samples in Figure 4(b) becomes more pronounced. Conversely, Figure 4(a) shows that after the expansion of distances between samples by LoRe, the gap between positive and negative samples is significantly smaller than the results in Figure 4(b). Another intriguing observation is that before projection, negative samples cluster in a narrow region, whereas positive samples distribute more uniformly across the space. It implies they occupy a narrow conical area in the high-dimensional space. Through LoRe, this conical shape can become more “flattened”. The observations indicate that LoRe enhances model performance by modulating the rate of distance diffusion among samples.
@@ 5 Conclusion
Acknowledgments
This work is supported in part by National Key R&D Program of China under Grant No. 2020AAA0109700, Guangdong Outstanding Youth Fund (Grant No. 2021B1515020061), National Natural Science Foundation of China (NSFC)under Grant No.61976233, Mobility Grant Award under Grant No. M-0461, Shenzhen Science and Technology Program (Grant No. GJHZ20220913142600001), Nansha Key RD Program under Grant No.2022ZD014, National Natural Science Foundation of China under Grant No.62006255, CAAl-Huawei MindSpore Open Fund. We thank MindSpore [^1] https://www.mindspore.cn/ for the partial support of this work, which is a new deep learning computing framwork.